OpenStudy (anonymous):

A compound is found to be 3 percent hydrogen, 32 percent phosphorus, and 65 percent oxygen. What is the empirical formula of this compound?

OpenStudy (anonymous):

First assume you have 100 grams of compound. 3% hydrogen=3/1=3 moles 32% phosphorous=32/31=1.03 moles 65% oxygen=65/16=4.0625 So ratio is 3:1:4 H=3 P=1 O=4 Now we will add these numbers as empirical formula So H3PO4 is empirical formula..

OpenStudy (anonymous):

I divided them by their molecular masses..

OpenStudy (chmvijay):

@Ryaan good work up but they are not molecular mass they are atomic mass :)